# DeformTime: Capturing Variable Dependencies with Deformable Attention for Time Series Forecasting

### v2

@@ D.7 Should MTS forecasting models mix variables in the embedding space?
The paper that presented the PatchTST method [32] provided some empirical evidence for the potential benefits of not modelling inter-variable dependencies. Contrary to that, our results demonstrate that this may not always hold for better models as well as different time series tasks and data sets. Based on our ILI rate forecasting experiments, we deduce that models that mix variables in the embedding space (Crossformer, LightTS, TimeMixer, iTransformer, and DeformTime) outperform the rest. Hence, and perhaps as expected, modelling inter-variable relationships should still be considered a viable approach in time series forecasting.
Appendix E Limitations
In this section, we summarise the potential limitations of this work. As with any machine learning model, the predictive capacity of a forecasting regression function is predominantly determined based on empirical evaluations. To this end, we have conducted extensive experiments. Nevertheless, there are undoubtedly more forecasting tasks that one could use to assess the performance of DeformTime. We have argued about the existence of potential issues within the currently deployed data sets in the evaluation of deep learning forecasters and we have proposed the use of new ones in this paper (alongside some of the previously established benchmarks). Although we hope that we will be able to share this data set with the broader community, at the time of submission we could only point to the application form for gaining access to the Google Health Trends API that we obtained the data from.
From a qualitative analysis standpoint, it would be interesting to provide further insights about DeformTime’s performance under a controlled amount of exogenous predictors. We do make the claim that DeformTime is designed to handle exogenous predictors better and experiment with new data sets that contain significantly more exogenous predictors compared to the established benchmark data sets (and show increased performance gains). However, we do not assess this under a more rigorous experimental design. We could add that our ablation study offers some level of interpretation for the contributions of different components. Nonetheless, further experiments would be required to more concretely understand DeformTime’s modelling strengths and weaknesses.
From a methodological perspective, there might be ways to further improve DeformTime. These, of course, could be considered as future avenues of research (as opposed to limitations). First, the deformations we are performing are localised. Perhaps a more expansive deformation approach would be capable of capturing more intricate relationships that a correlation-based neighbourhood omits. In addition, and this applies to all deep learning MTS models, DeformTime does not offer any confidence bounds for its forecasting estimates. We could arguably achieve that via conformal prediction or other approaches. However, how would that affect the accuracy of our mean estimates and the overall model complexity? This is certainly an avenue for follow up work, especially if a forecasting model is to be deployed in practical applications.
@@ Appendix G Supplementary ILI rate forecasting figures
Figure G3: 7 days ahead forecasts for all influenza seasons and models for England (ILI-ENG).
Figure G4: 14 days ahead forecasts for all influenza seasons and models for England (ILI-ENG).
Figure G5: 21 days ahead forecasts for all influenza seasons and models for England (ILI-ENG).
Figure G6: 28 days ahead forecasts for all influenza seasons and models for England (ILI-ENG).
Figure G7: 7 days ahead forecasts for all influenza seasons and models for US Region 2 (ILI-US2).
Figure G8: 14 days ahead forecasts for all influenza seasons and models for US Region 2 (ILI-US2).
Figure G9: 21 days ahead forecasts for all influenza seasons and models for US Region 2 (ILI-US2).
Figure G10: 28 days ahead forecasts for all influenza seasons and models for US Region 2 (ILI-US2).
Figure G11: 7 days ahead forecasts for all influenza seasons and models for US Region 9 (ILI-US9).
Figure G12: 14 days ahead forecasts for all influenza seasons and models for US Region 9 (ILI-US9).
Figure G13: 21 days ahead forecasts for all influenza seasons and models for US Region 9 (ILI-US9).
Figure G14: 28 days ahead forecasts for all influenza seasons and models for US Region 9 (ILI-US9).

### v3

@@ E.9 Should forecasting models mix variables in the embedding space?
The paper that presented the PatchTST method (Nie et al., 2023) provided some empirical evidence for the potential benefits of not modelling inter-variable dependencies. Contrary to that, our results demonstrate that this may not always hold for better-performing models as well as different time series forecasting tasks. Based on our ILI rate forecasting experiments, we deduce that models that mix variables in the embedding space while preserving the temporal structure (ModernTCN, Crossformer, LightTS, TimeMixer, and DeformTime) outperform the rest. Hence, and perhaps as expected, modelling inter-variable relationships should still be considered a viable (and arguably essential) approach in time series forecasting.
@@ Appendix G Supplementary ILI rate forecasting figures
Figure S2: 7 days ahead forecasts for all influenza seasons and models for England (ILI-ENG).
Figure S3: 14 days ahead forecasts for all influenza seasons and models for England (ILI-ENG).
Figure S4: 21 days ahead forecasts for all influenza seasons and models for England (ILI-ENG).
Figure S5: 28 days ahead forecasts for all influenza seasons and models for England (ILI-ENG).
Figure S6: 7 days ahead forecasts for all influenza seasons and models for US Region 2 (ILI-US2).
Figure S7: 14 days ahead forecasts for all influenza seasons and models for US Region 2 (ILI-US2).
Figure S8: 21 days ahead forecasts for all influenza seasons and models for US Region 2 (ILI-US2).
Figure S9: 28 days ahead forecasts for all influenza seasons and models for US Region 2 (ILI-US2).
Figure S10: 7 days ahead forecasts for all influenza seasons and models for US Region 9 (ILI-US9).
Figure S11: 14 days ahead forecasts for all influenza seasons and models for US Region 9 (ILI-US9).
Figure S12: 21 days ahead forecasts for all influenza seasons and models for US Region 9 (ILI-US9).
Figure S13: 28 days ahead forecasts for all influenza seasons and models for US Region 9 (ILI-US9).